40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 1949     Entry time: Thu Oct 12 16:57:51 2017
Author: Craig, awade 
Type: DailyProgress 
Category: BEAT 
Subject: Transmission Beatnote Recovered 

Just a quick update.  We recovered our beatnote and (maybe?) lowered ringing in our FSS boxes with some specific gain values.

South Laser Slow Control Voltage: -6.13 V

North Laser Slow Control Voltage: 3.54 V

North Cavity Temperature Control Power Supply Voltage: 11.25 V  (This swings our beatnote by about 250 MHz/V.  Also this setup is terrible, we need to remove all clipdoodles from the main setup.)

Slow PID Loop Time Step: 0.1 seconds

South Common Gain Slider: -3.5

South Fast Gain Slider: 2.28 (This one is very sensitive... moving it slightly causes ringing)

North Common Gain Slider: -3.13

North Fast Gain Slider: -1.29

The peak is still moving since we recently changed the North Cavity Temperature by 0.75 volts to put the locked region of the two cavities within 125 MHz of each other.  The cavity has a time constant of ~1 hour... we can and should measure this sometime.

Next step is to watch the slow control voltages over a long period of time along with the vac can temperature sensors to see if vac can temperature flucutations are still affecting our lock to the PLL.

Attachment 1: October12Beatnote.jpg  1.631 MB  | Hide | Hide all
October12Beatnote.jpg
ELOG V3.1.3-